### v1

@@ Abstract
The integration of Large Language Models (LLMs) with Knowledge Representation Learning (KRL) signifies a pivotal advancement in the field of artificial intelligence, enhancing the ability to capture and utilize complex knowledge structures. This synergy leverages the advanced linguistic and contextual understanding capabilities of LLMs to improve the accuracy, adaptability, and efficacy of KRL, thereby expanding its applications and potential. Despite the increasing volume of research focused on embedding LLMs within the domain of knowledge representation, a thorough review that examines the fundamental components and processes of these enhanced models is conspicuously absent. Our survey addresses this by categorizing these models based on three distinct Transformer architectures, and by analyzing experimental data from various KRL downstream tasks to evaluate the strengths and weaknesses of each approach. Finally, we identify and explore potential future research directions in this emerging yet underexplored domain, proposing pathways for continued progress.
@@
Knowledge graph, Large language model, Representation learning
@@ 1 Introduction
Large language models (LLMs) (e.g., BERT 18 , LLaMA 59 ) which represents a direction of ever-increasing model sizes pre-trained on larger corpora, have demonstrated powerful capabilities in solving natural language processing (NLP) tasks, including question answering 99 , text generation 100 and document understanding 101 . There are no clear and static thresholds regarding the model sizes. Early LLMs (e.g., BERT, RoBERTa) adopt an encoder architecture and show capabilities in text representation learning and natural language understanding. In recent years, more focus has been given to larger encoder-decoder 102 or decoder-only 103 architectures. As the model size scales up, such LLMs have also shown reasoning ability and even more advanced emergent ability 104 , exposing a strong potential for Artificial General Intelligence (AGI).
This inflection point, with the arrival of LLMs, marks a paradigm shift from explicit knowledge representation to a renewed focus on the hybrid representation of both explicit knowledge and parametric knowledge. As a popular approach for explicit knowledge representation, KGs are now widely investigated for the combination with Transformer-based LLMs, including pre-trained masked language models (PLMs) like BERT and RoBERTa, and more recent generative LLMs like the GPT series and LLaMA. Some works use LLMs to augment knowledge graph representation learning. In this survey, considering three directions, encoder-based methods, encoder-decoder-based method, and decoder-based methods. We present a better understanding of the shift from explicit knowledge representation to a renewed focus on the hybrid representation of both explicit knowledge and parametric knowledge.
@@
Cao et al. 22 and Biswas et al. 40 discuss recent advancements in knowledge graph representation learning, yet they inadequately address aspects related to the integration with large models. Pan et al. 42 and Pan et al. 43 explore the combination of knowledge graphs and large models, specifically addressing LLM4KG and KG4LLM; however, they provide limited coverage of representation learning. Consequently, there is currently no comprehensive review article dedicated to outlining the latest developments in the field of knowledge graph representation learning.
Contributions.   The notable contributions of this survey are summarized as follows:
  • Categorization of KRL. We systematically summarize the knowledge representation learning where large language models can be adopted into: encoder-based, encoder-decoder-based, and decoder-based methods.
  • Systematic Review of Techniques. We provide the most comprehensive overview of large language models on knowledge graph representation learning techniques. For different methods, we summarize the representative models, provide detailed illustrations, and make necessary comparisons.
  • Future Directions. We delve into the foundational principles of language models on graphs and propose six prospective avenues for future exploration.
@@
Organization.   The structure of this survey is organized as follows. Section 2 introduce the foundation by introducing the foundational concepts pertinent to this review, including Knowledge Graphs, Large Language Models, and Text-based Representation Learning. Sections 3 investigates the core methodologies employed in current research, categorizing them into three types of Transformer, respectively. Each of these sections further explores various sub-methodologies. In Section 4, we discuss the advantages of each type of approach from the performance evaluation results of each downstream task. Section 5 outlines potential future research directions that align with the discussed categories, proposing areas that promise significant advancements. The conclusion in Section 6 synthesizes the insights drawn from the review and highlights the implications for future research.
@@ 2.2 Large Language Model
Early pretraining methods in NLP, such as the Neural Language Model (NLM) and Word2vec, were static, meaning they could not adapt effectively to different semantic environments. Static models learn fixed word representations, which limits their ability to handle context variations. To overcome this limitation, dynamic pretraining techniques were developed, leading to Pretrained Language Model (PLM) like BERT (Bidirectional Encoder Representations from Transformers) and XLNet. These models dynamically adjust word representations based on context, significantly enhancing their adaptability and performance across various tasks.
@@
where $Q,K,V\in\mathds{R}^{N_{S}\times d_{k}}$ are the query, key, and value vectors for each word in the sentence, respectively. Following BERT, other masked language models are proposed, such as RoBERTa, ALBERT, and ELECTRA, with similar architectures and objectives of text representation.
Although the original Transformer paper was experimented on machine translation, it was not until the release of GPT-2 that language generation (aka. causal language modeling) became impactful on downstream tasks. Causal language modeling is the task of predicting the next word given the previous words in a sentence. The objective of causal language modeling is defined as:
@@
A prime example of the application of LLM in NLP is ChatGPT, which is fine-tuned from the generative pretrained transformer GPT-3.5. GPT-3.5 was trained on a diverse blend of text and code, making it adept at a wide range of tasks from conversation to code generation. ChatGPT leverages reinforcement learning from human feedback (RLHF), which has emerged as a promising approach to align large language models with human intent. This alignment is achieved through iterative tuning based on human feedback, ensuring the responses are more relevant and accurate.
@@
We would like to point out that the word “large” in LLM is not associated with a clear and static threshold to divide language models. “Large” actually refers to a direction in which language models are inevitably evolving, and larger foundational models tend to possess significantly more representation and generalization power. Hence, we define LLMs to encompass both medium-scale PLMs, such as BERT, and large-scale LMs, like GPT-4, as suggested.
@@ 2.3 Knowledge Representation Learning
Convolutional neural networks (CNNs) are effective for encoding entity descriptions and textual mentions, learning deep expressive features from textual information. DKRL(CNN) assumes that word order in entity descriptions involves implicit relations between entities that KGs may omit. It uses a five-layer CNN to discover these implicit relations. The vector embeddings of words, excluding stop words, are input to the CNN, with max-pooling and mean-pooling operations used to manage parameter space and filter noise. The non-linear output layer constructs the text-based entity representation.
@@
Models Structual embedding Textual embedding Extended model Combination mechanism WVV $r_{s}\in\mathbb{R}^{d}$ $h^{(r_{s})}_{d},t^{(r_{s})}_{d}\in\mathbb{R}^{d}$ SE /TransE/ $f_{r}(h,t)=-\|h^{(r_{s})}_{d}-r_{s}+t^{(r_{s})}_{d}\|$ DKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE RTKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $r_{d}\in\mathbb{R}^{d}$ TransE $f_{r}(h,t)=f_{s}+f_{d}$ RDRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE STKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE Joint $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE $h=g_{h}\otimes h_{s}+(1-g_{h})\otimes h_{d}$ $t=g_{t}\otimes t_{s}+(1-g_{t})\otimes t_{d}$ $r=r_{s}$ EDGE $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
@@
[…]
@@
Model Type Base Model Source Code 2019 ERNIE Encoder BERT https://github.com/thunlp/ERNIE KnowBert Encoder BERT https://github.com/allenai/kb KG-BERT Encoder BERT https://github.com/yao8839836/kg-bert 2020 MTL-KGC Encoder BERT https://github.com/bosung/MTL-KGC Pretrain-KGE Encoder BERT - K-BERT Encoder BERT https://github.com/autoliuweijie/K-BERT 2021 StAR Encoder BERT, RoBERTa https://github.com/wangbo9719/StAR_KGC KG-GPT2 Decoder GPT2 - MEM-KGC Encoder BERT - LaSS Encoder BERT, RoBERTa https://github.com/jhshen95/LASS KEPLER Encoder RoBERTa https://github.com/THU-KEG/KEPLER BLP Encoder BERT https://github.com/dfdazac/blp SimKGC Encoder BERT https://github.com/intfloat/SimKGC 2022 LP-BERT Encoder BERT, RoBERTa - PKGC Encoder BERT, RoBERTa, LUKE https://github.com/THU-KEG/PKGC Nayyeri et al. Encoder BERT https://github.com/ZihaoWang/Hypercomplex-KG-Embedding CoDEx Encoder Sentence-BERT - KGT5 Encoder-Decoder T5 https://github.com/apoorvumang/kgt5 OpenWorld KGC Encoder BERT - Lovelace et al. Encoder BERT, PubMedBERT https://github.com/justinlovelace/LM-KG-Completion kNN-KRL Encoder BERT https://github.com/zjunlp/KNN-KG LMKE Encoder BERT https://github.com/Neph0s/LMKE GenKGC Encoder-Decoder BART https://github.com/zjunlp/PromptKG/tree/main/research/GenKGC KG-S2S Encoder-Decoder T5 https://github.com/chenchens190009/KG-S2S 2023 LambdaKG Encoder-Decoder BERT, BART, T5 https://github.com/zjunlp/PromptKG/tree/main/lambdaKG CSPromp-KG Encoder BERT https://github.com/chenchens190009/CSProm-KG ReSKGC Encoder-Decoder T5 - StructGPT Decoder GPT 3.5 https://github.com/RUCAIBox/StructGPT ReasoningLM Encoder RoBERTa https://github.com/RUCAIBox/ReasoningLM KG-GPT Decoder GPT 3.5 https://github.com/jiho283/KG-GPT Keqing Decoder GPT 3.5, LLaMA 2 - AutoKG Decoder GPT 3.5 https://github.com/zjunlp/AutoKG KG-LLM Decoder ChatGLM, LLaMA 2 https://github.com/yao8839836/kg-llm KoPA Decoder Alpaca, LLaMA, GPT 3.5 https://github.com/zjukg/KoPA 2024 CD Decoder PaLM2 https://github.com/DavidLi0406/Contextulization-Distillation CP-KGC Decoder Qwen, LLaMA 2, GPT 4 https://github.com/sjlmg/CP-KGC KICGPT Decoder GPT 3.5 - KnowledgeNavigator Decoder GPT 3.5, LLaMA 2 -
Figure 2: A taxonomy of LLM enhanced KRL.
@@ 3 LLM Enhanced KRL Methods
In this section, we delve into the advanced methodologies for enhancing Knowledge Representation Learning using large language models. The section is structured into three main sections, each focusing on a different category of LLM-based methods. Section 3.1 explores Encoder-based methods, which utilize powerful encoders to generate dense vector representations of knowledge entities. Section 3.2 discusses Encoder-Decoder-based methods that leverage both encoding and decoding processes to capture and generate complex relations within knowledge graphs. Finally, Section 3.3 examines Decoder-based methods, emphasizing the generation and prediction capabilities of decoders to infer and expand knowledge representations.
@@ 3.1 Encoder-Based Methods
These works employ the encoder of Transformer to transform knowledge graph elements into textual sequences that are then processed to evaluate triple plausibility or to complete knowledge graphs. The contextual capabilities are central to assessing relations and entity attributes in the graph structure.
3.1.1 Multi-Level Encoder
The typical approach is illustrated in Fig. 4. This category primarily focuses on employing BERT or similar encoder-based pre-trained language models to convert knowledge graph elements into textual sequences, which are then used to assess the plausibility of triples. The common method is to use BERT deep contextual embeddings to enhance the representation of entities and relations, thereby improving accuracy in predicting plausible triples.
Figure 4: An overview of multi-level encoder method.
$\text{Input}=[CLS]+\text{Tokenize}(h)+[SEP]+\text{Tokenize}(r)+[SEP]+\text{Tokenize}(t)+[SEP]$ (4)
where [CLS] and [SEP] are special tokens used by BERT, and $\text{Tokenize}(\cdot)$ represents the tokenization process.
The classification task is to determine if the concatenated sequence correctly represents a valid triple from the KG. A logistic sigmoid function computes the probability of the triple being valid:
$s_{\tau}=\sigma(\mathbf{C}\mathbf{W}^{T})$ (5)
Here, $\mathbf{C}$ is the final hidden state of the [CLS] token, $\mathbf{W}$ is a learnable parameter matrix, and $\sigma$ denotes the sigmoid function.
The training uses a cross-entropy loss function to handle both positive and negative examples of triples:
$L=-\sum_{\tau\in D^{+}\cup D^{-}}\left(y_{\tau}\log(s_{\tau,0})+(1-y_{\tau})\log(s_{\tau,1})\right)$ (6)

### v2

@@ Abstract
The integration of Large Language Models (LLM) with Knowledge Representation Learning (KRL) signifies a significant advancement in the field of artificial intelligence (AI), enhancing the ability to capture and utilize both structure and textual information. Despite the increasing research on enhancing KRL with LLMs, a thorough survey that analyse processes of these enhanced models is conspicuously absent. Our survey addresses this by categorizing these models based on three distinct Transformer architectures, and by analyzing experimental data from various KRL downstream tasks to evaluate the strengths and weaknesses of each approach. Finally, we identify and explore potential future research directions in this emerging yet underexplored domain.
@@
Knowledge graph, Large language model, Knowledge representation learning
@@ 1 Introduction
Large language models (LLM) such as BERT [1], BART [2], and LLaMA [119], represent a trend towards increasingly larger models pre-trained on vast corpora. These models have exhibited remarkable capabilities in addressing various natural language processing (NLP) tasks, including text generation [3], sentiment analysis [4], and question answering [5]. There are no definitive and static thresholds for model sizes. Early LLM, like BERT and RoBERTa [6], utilize an encoder architecture and demonstrate proficiency in text representation learning and natural language understanding. More recently, there has been a shift towards larger encoder-decoder architectures or decoder architectures. As the size of these models grows, they exhibit advanced reasoning abilities and complex task-solving capabilities [7], indicating significant potential for a wide range of AI applications.
This inflection point, catalyzed by the introduction of LLM, heralds a paradigm shift from structured knowledge representation to the integration of structured and textual representations. Knowledge Graphs (KG), once primarily used for structured knowledge representation, are now being extensively explored in conjunction with Transformer-based LLM, such as pre-trained language models (PLMs) and recent generative LLM. Various studies leverage LLM to refine the processes of knowledge representation learning (KRL). In this survey, we categorize the approaches into three distinct methodologies: encoder-based, encoder-decoder-based, and decoder-based methods. This categorization aids in understanding the shift from purely structured representation to a hybrid model that effectively combines structured and textual knowledge representations.
@@
There are few existing surveys that investigate the enhancement of KRL by LLM. Cao et al. [8] and Biswas et al. [9] discuss recent advancements in KRL, yet they inadequately address aspects related to the integration with LLM. Pan et al. [10] and Pan et al. [11] explore the combination of KG and LLM, specifically addressing LLM4KG and KG4LLM. However, they provide limited coverage of KRL. Consequently, there is currently no comprehensive review article dedicated to outlining the latest developments in the field of KRL.
Contributions. The contributions of this survey can be summarized as follows:
  • LLM Enhanced KRL Survey. This work provides the first comprehensive survey on the LLM enhanced KRL. Specifically, we analyze how different encoder-based, encoder-decoder-based, and decoder-based methods have been enhanced through the integration with LLM, offering a new dimension to understanding their capabilities and limitations in KRL.
  • Categorization of KRL. We categorize the existing methods in KRL according to their foundational LLM architectures, providing detailed descriptions and comparisons of these methods from the perspective of their computational models and the enhancement strategies employed.
  • Future Directions. We propose some directions for future research in KRL by highlighting unique properties of LLM enhanced methods. These insights are expected to inspire advancements not only within the realm of KRL but also in broader fields like NLP, leveraging the robust capabilities of LLM.
@@
Organization. The structure of this survey is organized as follows. Section 2 (§2) introduces the foundational concepts that are essential to this survey, including Knowledge Graph, Large Language Model, and Knowledge Representation Learning. Sections 3 (§3) investigates the core methodologies employed in current research, categorizing them into three types of Transformer, respectively. Each of these sections further explores various sub-methodologies. In Section 4 (§4), we discuss the advantages of each type of approach from the performance evaluation results of each downstream task. Section 5 (§5) outlines potential future directions that align with the discussed categories. The conclusion in Section 6 (§6) provides a comprehensive summary of the survey.
@@ 2.2 Large Language Model
Early pretraining methods in NLP, such as the Neural Language Model (NLM) [12] and Word2vec [13], were static, meaning they could not adapt effectively to different semantic environments. Static models learn fixed word representations, which limits their ability to handle context variations. To overcome this limitation, dynamic pretraining techniques were developed, leading to Pretrained Language Model (PLM) like BERT (Bidirectional Encoder Representations from Transformers) and XLNet [14]. These models dynamically adjust word representations based on context, significantly enhancing their adaptability and performance across various tasks.
@@
where $Q,K,V\in\mathds{R}^{N_{S}\times d_{k}}$ are the query, key, and value vectors for each word in the sentence, respectively. Following BERT, other masked language models are proposed, such as RoBERTa, ALBERT [15], and ELECTRA [16], with similar architectures and objectives of text representation.
Although the original Transformer paper was experimented on machine translation, it was not until the release of GPT-2 [17] that language generation (aka. causal language modeling) became impactful on downstream tasks. Causal language modeling is the task of predicting the next word given the previous words in a sentence. The objective of causal language modeling is defined as:
@@
A prime example of the application of LLM in NLP is ChatGPT, which is fine-tuned from the generative pretrained transformer GPT-3.5. GPT-3.5 was trained on a diverse blend of text and code, making it adept at a wide range of tasks from conversation to code generation. ChatGPT leverages reinforcement learning from human feedback (RLHF) [18], which has emerged as a promising approach to align large language models with human intent. This alignment is achieved through iterative tuning based on human feedback, ensuring the responses are more relevant and accurate.
@@
We would like to point out that the word ‘large’ in LLM is not associated with a clear and static threshold to divide language models. ‘Large’ actually refers to a direction in which language models are inevitably evolving, and larger foundational models tend to possess significantly more representation and generalization power. Hence, we define LLM to encompass both medium-scale PLMs, such as BERT, and large-scale LMs, like GPT-4, as suggested.
@@ 2.3 Knowledge Representation Learning
Convolutional neural networks (CNNs) are effective for encoding entity descriptions and textual mentions, learning deep expressive features from textual information. DKRL(CNN) assumes that word order in entity descriptions involves implicit relations between entities that KG may omit. It uses a five-layer CNN to discover these implicit relations. The vector embeddings of words, excluding stop words, are input to the CNN, with max-pooling and mean-pooling operations used to manage parameter space and filter noise. The non-linear output layer constructs the text-based entity representation.
@@
Models Structual embedding Textual embedding Extended model Combination mechanism
WVV $r_{s}\in\mathbb{R}^{d}$ $h^{(r_{s})}_{d},t^{(r_{s})}_{d}\in\mathbb{R}^{d}$ SE /TransE/ $f_{r}(h,t)=-\|h^{(r_{s})}_{d}-r_{s}+t^{(r_{s})}_{d}\|$
DKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
RTKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $r_{d}\in\mathbb{R}^{d}$ TransE $f_{r}(h,t)=f_{s}+f_{d}$
RDRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
STKRL $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
Joint $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE $h=g_{h}\otimes h_{s}+(1-g_{h})\otimes h_{d}$
$t=g_{t}\otimes t_{s}+(1-g_{t})\otimes t_{d}$
$r=r_{s}$
EDGE $h_{s},t_{s},r_{s}\in\mathbb{R}^{d}$ $h_{d},t_{d}\in\mathbb{R}^{d}$ TransE
@@
Notation Description
$h$, $r$, $t$ Head entity, Relation, Tail entity
$\mathbf{X}{(h,r)}$, $\mathbf{X}(t)$ Input sequences for head-relation and tail entity
$D$, $D^{+}$, $D^{-}$ Set of all triples, positive triples, negative triples
$\mathbf{u}$, $\mathbf{v}$ Contextualized embeddings of head-relation pair and tail entity
$s_{c}$, $s_{d}$ Classification score, Distance score
$\mathbf{h}$, $\mathbf{r}$, $\mathbf{t}$ Embeddings for head entity, relation, and tail entity
$L_{c}$, $L_{d}$, $L$ Classification loss, Distance-based loss, Total loss
$\lambda$, $\gamma$, $n$, $K$ Margin, Weight factor, Number of negative samples, Sum limit
$\sigma$ Sigmoid function
$\alpha$, $\beta$ Auxiliary loss weights
$\mathbf{K}$, $I$, $X$ Transformed structural embeddings, Instruction prompt, Triple prompt
@@
Year Model Type Base Model Source Code
2019 KG-BERT Encoder BERT Github
2020 MTL-KGC Encoder BERT Github
Pretrain-KGE Encoder BERT -
K-BERT Encoder BERT Github
2021 StAR Encoder BERT, RoBERTa Github
KG-GPT2 Decoder GPT2 -
BERT-ResNet Encoder BERT Github
MEM-KGC Encoder BERT -
LaSS Encoder BERT, RoBERTa Github
KEPLER Encoder RoBERTa Github
BLP Encoder BERT Github
SimKGC Encoder BERT Github
MLMLM Encoder RoBERTa Github
2022 LP-BERT Encoder BERT, RoBERTa -
PKGC Encoder BERT, RoBERTa, LUKE Github
KGT5 Encoder-Decoder T5 Github
OpenWorld KGC Encoder BERT -
Lovelace et al. Encoder BERT, PubMedBERT Github
$k$NN-KGE Encoder BERT Github
LMKE Encoder BERT Github
GenKGC Encoder-Decoder BART Github
KG-S2S Encoder-Decoder T5 Github
2023 LambdaKG Encoder-Decoder BERT, BART, T5 Github
CSPromp-KG Encoder BERT Github
ReSKGC Encoder-Decoder T5 -
ReasoningLM Encoder RoBERTa Github
KG-LLM Decoder ChatGLM, LLaMA 2 Github
KoPA Decoder Alpaca, LLaMA, GPT 3.5 Github
2024 CD Decoder PaLM2 Github
CP-KGC Decoder Qwen, LLaMA 2, GPT 4 Github
KICGPT Decoder GPT 3.5 -
@@ 3 LLM Enhanced KRL Methods
LLM have significantly advanced the field of KRL. These models leverage vast amounts of data and sophisticated architectures to generate context-aware representations of knowledge. This section explores various methods for enhancing KRL using LLM, categorized into three main approaches: encoder-based methods (§3.1), encoder-decoder-based methods (§3.2), and decoder-based methods (§3.3). Each category leverages different aspects of LLM to optimize the encoding, decoding, and overall integration of knowledge graph data, ultimately improving the accuracy and utility of knowledge representations in diverse applications. The evolution of these models is illustrated in Fig. 2.
@@ 3.1 Encoder-Based Methods
Encoder-based methods primarily utilize encoders such as BERT [1] and RoBERTa [6]. These methods are distinguished based on how they encode KG triples. In triple encoding (§3.1.1), the head entity, relation, and tail entity are encoded together as a single unit. Translation encoding (§3.1.2), on the other hand, encodes the head entity and relation together, while the tail entity is encoded separately, with optimization performed using vector space distances. Independent encoding (§3.1.3) treats the head entity, relation, and tail entity as separate inputs, encoding each independently.
3.1.1 Triple Encoding
Triple encoding methods aim to represent the entire KG triple as a single unit. This approach leverages the combined semantic and relational information of the head entity, relation, and tail entity, providing a holistic representation that can enhance downstream tasks. An example of a typical model KG-BERT [19], that using triple encoding is illustrated in Fig. 3.
Figure 3: An overview of triple encoding method.
It treats triples as textual sequences and fine-tunes a pre-trained BERT model to predict the plausibility of a given triple. The model takes as input the descriptions of the head entity $h$, relation $r$, and tail entity $t$:
[…]
where $\text{Tok}_{hi}$, $\text{Tok}_{ri}$, and $\text{Tok}_{ti}$ represent the tokens of the head entity, relation, and tail entity descriptions, respectively. The model computes the final hidden vector of the special [CLS] token, which serves as the aggregate sequence representation. The plausibility score of a triple $\tau=(h,r,t)$ is then computed as follows:
$s_{\tau}=\text{sigmoid}(CW^{T})$ (5)
where $C$ is the final hidden state corresponding to the [CLS] token, and $W$ is the classification layer weight. The loss function for training is the cross-entropy loss computed over the positive and negative triples:
$L=-\sum_{\tau\in D^{+}\cup D^{-}}(y_{\tau}\log(s_{\tau 0})+(1-y_{\tau})\log(s_{\tau 1}))$ (6)